OpenStudy (anonymous):

The length of a rectangle is 3b + 2 and the width is 9b + 6. What is the ratio of its length to its width? Simplify your answer.

OpenStudy (anonymous):

\[\frac{ 3b +2 }{ 9b + 6 }\] \[\frac{ 3b +2 }{ 3(3b + 2) }\] \[\frac{ 1 }{ 3 } * \frac{ 3b +2 }{ (3b + 2) }\] \[\frac{ 1 }{ 3 } * 1\] \[\frac{ 1 }{ 3 }\]